18. One of three possible hypotheses H,, H2, or H, can explain the occurrence of a road accident. The probability of the accident if H is correct is .75. The probability of the accident if H2 is correct is .80. The probability of the accident if H is correct is .85. H has probability 4 of occurring, H2 has probability .5 of occurring, and H has probability 1 of occurring. De termine the most probable hypothesis to explain an observed accident
Answer #1

Let p1 = Probability of hypothesis H1 explaining an accident correctly when the accident occurs

p2 = Probability of hypothesis H2 explaining an accident correctly when the accident occurs

p3 = Probability of hypothesis H3 explaining an accident correctly when the accident occurs

p1 = P(H1 explaining correctly)*P(H1 occurs)

      = 0.75*0.4

      = 0.3

p2 = P(H2 explaining correctly)*P(H2 occurs)

      = 0.8*0.5

      = 0.4

p3 = P(H3 explaining correctly)*P(H3 occurs)

      = 0.85*0.1

      = 0.085

p2 is the highest among p1, p2 and p3

Hence the "most probable" hypothesis to explain an observed accident is H2